OpenStudy (anonymous):

3.) If a a circle has radius 7, what is the area? A = _____ 153.86 14.00 43.96 3.50

OpenStudy (anonymous):

a=Pi(r)^2 a=Pi(7)^2 a=153.86

OpenStudy (anonymous):

\[Area = pi*(radius^2)\]So around 153.